twice a number is 60 more than five times the number. what is the number?
Found 2 solutions by checkley77, jsmallt9:
Answer by checkley77(12844)

2X+60=5X
2X-5X=-60
-3X=-60
X=-60/-3
X=20 ANS.
PROOF:
2*20+60=5*20
40+60=100
100=100


Answer by jsmallt9(3758)

Let's call "a number" x. Then twice a number would be 2x. And 5 times the number is 5x. Now we can translate "twice a number is 60 more than five times the number":
2x = 60 + 5x
Now we solve this for x. Subtract 5x from each side:
-3x = 60
Divide both sides by -3:
x = -20

Let's check. Twice -20 is -40. 5 times -20 is -100. And -40 is 60 more than -100.
So the number is -20.

Note: a solution provided by another is incorrect. It uses an incorrect equation and then checks/proves the incorrect answer with the incorrect equation.